OpenStudy (anonymous):

1 Calculate the number of moles of : (a) 2.0 g calcium, Ca

OpenStudy (anonymous):

molecular mass of ca:40 so 40 gm of ca =1 mole 2gm of ca=1/40*2=1/20=0.05mole